OpenStudy (anonymous):

Find the coordinates of the midpoint of the segment whose endpoints are (10, 6) and (-4, 8).

OpenStudy (anonymous):

9,1 7,-1 3,7 >>>>>>>>>> ???????

OpenStudy (anonymous):

Just sum the x coordinates and then divide by 2.. Similarly add the y coordinates and then divide by 2..

OpenStudy (anonymous):

7,-1?

OpenStudy (anonymous):

\[Midpoint(x,y) = (\frac{x_1 + x_2}{2}, \frac{y_1 + y_2}{2})\]

OpenStudy (anonymous):

Check it once again..

OpenStudy (anonymous):

9,1?

OpenStudy (anonymous):

\[10 + (-4) \implies 10 - 4 = 6 \implies \frac{6}{2} = ??\]

OpenStudy (anonymous):

See in case of addition : +(-) = - So : \[\frac{10 + (-4)}{2} \implies \frac{10 - 4}{2} = ??\]

OpenStudy (anonymous):

i'm so confused. -_-

OpenStudy (anonymous):

Where?? Can you tell me?

OpenStudy (anonymous):

like I understand 6/2 but my schoolwork has none of those answers even if I divide.

OpenStudy (anonymous):

See if you have 10 and +4 as x coordinates you will add like this: \[10 + (+4) = 10 + 4 = 14\] In that case you will get 7, no doubt.. But here it is -4 not +4.

OpenStudy (anonymous):

What is the third answer choice??

OpenStudy (anonymous):

Simply tell me what is 6/2 = ??

OpenStudy (jim766):

here is what I do 10, 6 -4 , 8 Stack the points ________ 6 14 Add divide by 2 /2 /2 (3, 7)